# Boosting Latent Diffusion with Perceptual Objectives

### v1

@@ 1 Introduction
Latent diffusion models (LDMs) (Rombach et al., 2022) have enabled considerable advances in image generation, and elevated the problem of generative image modeling to a level where it has become available as a technology to the public. A critical part to this success is to define the generative model in the latent space of an autoencoder (AE), which reduces the resolution of the representation over which the model is defined, thereby making it possible to scale diffusion methods to larger datasets, resolutions, and architectures than original pixel-based diffusion models (Dhariwal and Nichol, 2021; Sohl-Dickstein et al., 2015).
@@
In this work, we propose to alleviate this autoencoder-diffusion disconnect and propose to include the AE decoder in the training objective of LDM. In particular, we introduce latent perceptual loss (LPL) that acts on the decoder’s intermediate features to enrich the training signal of LDM. This is similar to the use of perceptual losses for image-to-image translation tasks (Johnson et al., 2016; Zhang et al., 2018), but we apply this idea in the context of generative modeling and use the feature space of the pre-trained AE decoder rather than that of an external pre-trained discriminative network. Our latent perceptual loss results in sharper and more realistic images, and leads to better structural consistency than the baseline – see Figure 1. We validate LPL on three datasets of different sizes – the commonly used datasets ImageNet-1k (1M data points) and CC12M (12M data points), and additionally a private dataset S320M (320M data points) – as well as three generative models formulation – DDPM (Ho et al., 2020) with velocity and epsilon prediction, and conditional flow matching model (Lipman et al., 2023). In our experiments, we report standard image generative model metrics – such as FID (Heusel et al., 2017), CLIPScore (Hessel et al., 2021), as well as Precision and Recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019). Our experiments show that the use of LPL leads to consistent performance boosts between 6% and 20% in terms of FID. Our qualitative analysis further highlights the benefits of LPL, showing images that are sharp and contain high-frequency image details.
In summary, our contributions are:
@@ 2 Related work
Diffusion models. The generative modeling landscape has been significantly impacted by diffusion models, surpassing previous state-of-the-art GAN-based methods (Brock et al., 2019; Karras et al., 2019, 2020, 2021). Diffusion models offer advantages such as more stable training and better scalability, and were successfully applied to a wide range of applications, including image generation (Chen et al., 2024; Ho et al., 2020), video generation (Ho et al., 2022b; Singer et al., 2023), music generation (Levy et al., 2023; San Roman et al., 2023), and text generation (Wu et al., 2023). Various improvements of the framework have been proposed, including different schedulers (Lin et al., 2024; Hang and Gu, 2024), loss weights (Choi et al., 2022; Hang et al., 2023), and more recently generalizations of the framework with flow matching (Lipman et al., 2023). In our work we evaluate the use of our latent perceptual loss in three different training paradigms: DDPM under noise and velocity prediction, as well as flow-based training with the optimal transport path.
@@
Latent diffusion. Due to the iterative nature of the reverse diffusion process, training and sampling diffusion models is computationally demanding, in particular at high resolution. Different approaches have been explored to generate high-resolution content. For example, Ho et al. (2022a) used a cascaded approach to progressively add high-resolution details, by conditioning on previously generated lower resolution images. A more widely adopted approach is to define the generative model in the latent space induced by a pretrained autoencoder (Rombach et al., 2022), as previously explored for discrete autoregressive generative models (Esser et al., 2021). Different architectures have been explored to implement diffusion models in the latent space, including convolutional UNet-based architectures (Rombach et al., 2022; Podell et al., 2024), and more recently transformer-based ones (Peebles and Xie, 2023; Chen et al., 2024; Gao et al., 2023; Esser et al., 2024) which show better scaling performance. Working in a lower-resolution latent space accelerates training and inference, but training models using a loss defined in the latent space also deprives them from matching high-frequency details from the training data distribution. Earlier approaches to address this problem include the use of a refiner model (Podell et al., 2024), which consists of a second diffusion model trained on high-resolution high-quality data that is used to noise and denoise the initial latents, similar to how SDEdit works for image editing (Meng et al., 2022). Our latent perceptual loss addresses this issue in an orthogonal manner by introducing a loss defined across different layers of the AE decoder in the latter stages of the training process. Our approach avoids the necessity of training on specialized curated data (Dai et al., 2023), and does not increase the computational cost of inference.
@@
Perceptual losses. The use of internal features of a fixed, pre-trained deep neural network to compare images or image distributions has become common practice as they have been found to correlate to some extent with human judgement of similarity (Johnson et al., 2016; Zhang et al., 2018). An example of this is the widely used Fréchet Inception Distance (FID) to assess generative image models (Heusel et al., 2017). Such “perceptual” distances have also been found to be effective as a loss to train networks for image-to-image tasks and boost image quality as compared to using simple $\ell_{1}$ or $\ell_{2}$ reconstruction losses. They have been used to train autoencoders (Esser et al., 2021), models for semantic image synthesis (Isola et al., 2017; Berrada et al., 2024b) and super-resolution (Suvorov et al., 2022; Jo et al., 2020), and to assess the sample diversity of generative image models (Schönfeld et al., 2021; Astolfi et al., 2024). In addition, recent works propose variants that do not require pretrained image backbones (Amir and Weiss, 2021; Czolbe et al., 2020; Veeramacheneni et al., 2023). An et al. (2024); Song et al. (2023) employed LPIPS as metric function in pixel space to train cascaded diffusion models and consistency models, respectively. Most closely related to our work, Kang et al. (2024) used a perceptual loss defined over latents to distill LDMs to conditional GANs, but used a separate image classification network trained over latents rather than the autoencoder’s decoder to obtain the features for this loss. In summary, compared to prior work on perceptual losses, our work is different in that (i) LPL is defined over the features of the decoder – that maps from latent space to RGB pixel space, rather than using a network that takes RGB images as input – and (ii) we use LPL to train latent diffusion models.
@@ 3 Using the latent decoder to define a perceptual loss
3.1 Latent diffusion and the MSE objective
@@
We use $F_{\beta}$ to refer to an autoencoder that consists of two modules. The encoder, $F^{e}_{\beta}$, maps RGB images $x_{0}\in\mathbb{R}^{H\times W\times 3}$ to a latent representation $z_{0}\in\mathbb{R}^{H/d\times W/d\times C}$, where $d$ is the spatial downscaling factor, and $C$ the channel dimension of the autoencoder. The decoder, $F^{d}_{\beta}$, maps from the latent space to the RGB image space. In LDM, the diffusion model, $D_{\Theta}$, with parameters $\Theta$ is defined over the latent representation of the autoencoder. We follow a typical setting, see e.g. (Peebles and Xie, 2023; Chen et al., 2024; Rombach et al., 2022), where we use a fixed pre-trained autoencoder with a downsampling factor of $d=8$ and a channel capacity of $C=4$.
@@
Training Objective. The diffusion formulation results in an objective function that is a lower-bound on the log-likelihood of the data. In the DDPM paradigm (Ho et al., 2020), the variational lower bound can be expressed as a sum of denoising score matching losses (Vincent, 2011), and the objective function can be written as $\mathcal{L}=\sum_{t}\mathcal{L}_{t}$, where $\mathcal{L}_{t}=\mathcal{D}_{\textrm{KL}}\left[q\left(x_{t-1}|x_{t},x_{0}\right)||p_{\theta}\left(x_{t-1}|x_{t}\right)\right]=\mathbb{E}_{x_{0},\epsilon,t}\left[\frac{\beta_{t}}{{(1-\beta_{t})(1-\alpha_{t})}}\cdot\lVert\epsilon-\epsilon_{\theta}(x_{t},t)\rVert^{2}\right]$. Ho et al. (2020) observed that disregarding the time step specific weighting resulted in improved sample quality, and introduced a simplified noise reconstruction objective, known as epsilon prediction, where the objective is the average of the MSE loss between the predicted noise and the noise vector added to the image, $\mathcal{L}_{\text{simple}}=\sum_{t}\mathbb{E}_{x_{0},\epsilon,t}\left[\lVert\epsilon-\epsilon_{\theta}(x_{t},t)\rVert^{2}\right]$. The underlying idea is that the better the noise estimation, the better the final sample quality. An equivalent way to interpret this objective is through reparameterization of the target to the original latents, $\mathcal{L}_{\text{simple}}=\sum_{t}\lambda_{t}\cdot\mathbb{E}_{x_{0},\epsilon,t}\left[\lVert x_{0}-\hat{x}_{0}(x_{t},t;\theta)\rVert^{2}\right]$, where $\hat{x}_{0}(x_{t},t;\theta)=\sqrt{1+\sigma_{t}^{2}}x_{t}-\sigma_{t}\hat{\epsilon}_{t}$ and $\lambda_{t}=1/\sigma_{t}^{2}$.
@@
We note that the presence of $\ell_{2}$ in the LDM objective has some important implications. First, the $\ell_{2}$ norm treats all pixels in the latents as equally important and disregards the downstream structure induced by the decoder whose objective is to reconstruct the image from its latents. This is problematic because the autoencoder’s latent space has a highly irregular structure and is not equally influenced by the different pixels in the latent code. Thus, optimizing the $\ell_{2}$ distance in the diffusion model latent space could be different from optimizing the perceptual distance between images. Second, while an $\ell_{2}$ objective is theoretically justified in the original DDPM formulation, generative models trained with an $\ell_{2}$ reconstruction objective have been observed to produce blurry images, as is the case e.g. for VAE models (Kingma and Welling, 2014).
The problem of blurry images due to $\ell_{2}$ reconstruction losses has been addressed through the use of perceptual losses such as LPIPS (Zhang et al., 2018), which provide a significant boost to the image quality in settings such as autoencoding (Esser et al., 2021), super-resolution Ledig et al. (2017) and image-to-image generative models (Isola et al., 2017; Park et al., 2019). However, in the case of LDM, a perceptual loss cannot be used directly on the predicted latents, instead the latents would need to be decoded to an RGB image and then entered into a feature-extraction network – introducing significant overhead in terms of memory and computation. To avoid such overheads, we develop an alternative perceptual loss that operates directly on the feature space of the decoder.
@@ 3.2 Latent perceptual loss
We propose a perceptual loss that operates on the features of the latent-to-RGB decoder, $F^{d}_{\beta}$, extracted at different depths in the decoder. Let $z_{t}=a_{t}z_{0}+b_{t}\epsilon_{t}$ be a noisy sample in the diffusion model latent space at time $t$, and $\hat{z}_{0}=(z_{t}-b_{t}D(z_{t},\sigma_{t};\Theta))/a_{t}$ the corresponding estimated noise-free latent at time $t=0$. We compute two sets of hierarchies of $L$ decoder features, $\{\phi_{l}\}_{l=1}^{L}$, and, $\{\hat{\phi}_{l}\}_{l=1}^{L}$, by decoding both the original, $z_{0}$, and estimated latents, $\hat{z}_{0}$ (where for brevity we drop the dependence of $\hat{z}_{0}$ on $t$):
$\begin{cases}\phi_{1},...,\phi_{L}=\big{(}F_{\beta}^{d,l}(z_{0})\big{)}_{l\in[\![1,L]\!]},\\ \hat{\phi}_{1},...,\hat{\phi}_{L}=\big{(}F_{\beta}^{d,l}(\hat{z}_{0})\big{)}_{l\in[\![1,L]\!]}.\end{cases}$ (1)
@@
$\mathcal{L}_{\textrm{LPL}}=\mathbb{E}_{t\in\mathcal{T},\epsilon\sim\mathcal{N}(0,I),x_{0}\in D_{\mathcal{X}}}\left[\delta_{\sigma_{t}\leq\tau_{\sigma}}\sum_{l=1}^{L}\frac{\omega_{l}}{C_{l}}\sum_{c=1}^{C_{l}}\big{\lVert}\rho_{l,c}(\hat{\phi}_{l,c})\odot\big{(}\phi^{\prime}_{l,c}-\hat{\phi}^{\prime}_{l,c}\big{)}\big{\rVert}_{2}^{2}\right].$ (2)
where $\phi^{\prime}_{l}$ is the standardized version of $\phi_{l}$ across the channel dimension, $\rho_{l,c}(\hat{\phi}_{l,c})$ is a binary mask masking the detected outliers in the feature map $\hat{\phi}_{l,c}$, $\omega_{l}$ is a depth-specific weighting and $C_{l}$ the channel dimensionality of the feature tensor. Note that we better explain these terms later in this section. Moreover, to reduce both the LPL computational complexity and memory overhead, we only apply our loss for high signal-to-noise ratios (SNR). In particular, we impose a hard threshold $\tau_{\sigma}$ and only apply the loss if the SNR is higher than it $\delta_{\sigma_{t}\leq\tau_{\sigma}}(\sigma_{t})$.
The LPL loss is applied in conjunction with the standard diffusion loss, resulting in the following training objective.
$\mathcal{L}_{\textrm{tot}}=\mathcal{L}_{\textrm{Diff}}+w_{\textrm{LPL}}\cdot\mathcal{L}_{\textrm{LPL}}$ (3)
@@
Framework $a_{t}$ $b_{t}$ $\hat{x}_{0}$
DDPM-$\epsilon_{t}$ $1/\sqrt{1+\sigma_{t}^{2}}$ $\sigma_{t}/\sqrt{1+\sigma_{t}^{2}}$ $\sqrt{1+\sigma_{t}^{2}}x_{t}-\sigma_{t}D(x_{t},t;\Theta)$
DDPM-$v_{\theta}$ $1/\sqrt{1+\sigma_{t}^{2}}$ $\sigma_{t}/\sqrt{1+\sigma_{t}^{2}}$ $\frac{1}{\sqrt{\sigma_{t}^{2}+1}}\big{(}x_{t}-\sigma_{t}D(x_{t},t;\Theta)\big{)}$
Flow-OT $1-t$ $t$ $\big{(}x_{t}-tD(x_{t},t;\Theta)\big{)}/(1-t)$
Table 1: Summary of the formula for the estimate of the clean image corresponding to the different formulations. Using the following parameterization, $\forall t,x_{t}=a_{t}x_{0}+b_{t}\epsilon_{t}$.

### v2

@@ 1 Introduction
Latent diffusion models (LDMs) (Rombach et al., 2022) have enabled considerable advances in image generation, and elevated the problem of generative image modeling to a level where it has become available as a technology to the public. A critical part to this success is to define the generative model in the latent space of an autoencoder (AE), which reduces the resolution of the representation over which the model is defined, thereby making it possible to scale diffusion methods to larger datasets, resolutions, and architectures than original pixel-based diffusion models (Dhariwal & Nichol, 2021; Sohl-Dickstein et al., 2015).
@@
In this work, we propose to alleviate this autoencoder-diffusion disconnect and propose to include the AE decoder in the training objective of LDM. In particular, we introduce latent perceptual loss (LPL) that acts on the decoder’s intermediate features to enrich the training signal of LDM. This is similar to the use of perceptual losses for image-to-image translation tasks (Johnson et al., 2016; Zhang et al., 2018), but we apply this idea in the context of generative modeling and use the feature space of the pre-trained AE decoder rather than that of an external pre-trained discriminative network.
Our latent perceptual loss results in sharper and more realistic images, and leads to better structural consistency than the baseline – see Figure 1. We validate LPL on three datasets of different sizes – the commonly used datasets ImageNet-1k (1M data points) and CC12M (12M data points), and additionally a private dataset S320M (320M data points) – as well as three generative models formulation – DDPM (Ho et al., 2020) with velocity and epsilon prediction, and conditional flow matching model (Lipman et al., 2023). In our experiments, we report standard image generative model metrics – such as FID (Heusel et al., 2017), CLIPScore (Hessel et al., 2021), as well as Precision and Recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019). Our experiments show that the use of LPL leads to consistent performance boosts between 6% and 20% in terms of FID. Our qualitative analysis further highlights the benefits of LPL, showing images that are sharp and contain high-frequency image details. In summary, our contributions are:
  • We identify a slight disconnect between latent and pixel-space diffusion which can lead to suboptimal results when training latent diffusion and flow models.
@@ 2 Related work
Diffusion models. The generative modeling landscape has been significantly impacted by diffusion models, surpassing previous state-of-the-art GAN-based methods (Brock et al., 2019; Karras et al., 2019, 2020, 2021). Diffusion models offer advantages such as more stable training and better scalability, and were successfully applied to a wide range of applications, including image generation (Chen et al., 2024; Ho et al., 2020), video generation (Ho et al., 2022b; Singer et al., 2023), music generation (Levy et al., 2023; San Roman et al., 2023), and text generation (Wu et al., 2023). Various improvements of the framework have been proposed, including different schedulers (Lin et al., 2024; Hang & Gu, 2024), loss weights (Choi et al., 2022; Hang et al., 2023b), and more recently generalizations of the framework with flow matching (Lipman et al., 2023). In our work we evaluate the use of our latent perceptual loss in three different training paradigms: DDPM under noise and velocity prediction, as well as flow-based training with the optimal transport path.
@@
Latent diffusion. Due to the iterative nature of the reverse diffusion process, training and sampling diffusion models is computationally demanding, in particular at high resolution. Different approaches have been explored to generate high-resolution content. For example, Ho et al. (2022a) used a cascaded approach to progressively add high-resolution details, by conditioning on previously generated lower resolution images. A more widely adopted approach is to define the generative model in the latent space induced by a pretrained autoencoder (Rombach et al., 2022), as previously explored for discrete autoregressive generative models (Esser et al., 2021). Different architectures have been explored to implement diffusion models in the latent space, including convolutional UNet-based architectures (Rombach et al., 2022; Podell et al., 2024), and more recently transformer-based ones (Peebles & Xie, 2023; Chen et al., 2024; Gao et al., 2023; Esser et al., 2024) which show better scaling performance. Working in a lower-resolution latent space accelerates training and inference, but training models using a loss defined in the latent space also deprives them from matching high-frequency details from the training data distribution. Earlier approaches to address this problem include the use of a refiner model (Podell et al., 2024), which consists of a second diffusion model trained on high-resolution high-quality data that is used to noise and denoise the initial latents, similar to how SDEdit works for image editing (Meng et al., 2022). Our latent perceptual loss addresses this issue in an orthogonal manner by introducing a loss defined across different layers of the AE decoder in the latter stages of the training process. Our approach avoids the necessity of training on specialized curated data (Dai et al., 2023), and does not increase the computational cost of inference.
@@
Perceptual losses. The use of internal features of a fixed, pre-trained deep neural network to compare images or image distributions has become common practice as they have been found to correlate to some extent with human judgement of similarity (Johnson et al., 2016; Zhang et al., 2018). An example of this is the widely used Fréchet Inception Distance (FID) to assess generative image models (Heusel et al., 2017). Such “perceptual” distances have also been found to be effective as a loss to train networks for image-to-image tasks and boost image quality as compared to using simple $\ell_{1}$ or $\ell_{2}$ reconstruction losses. They have been used to train autoencoders (Esser et al., 2021), models for semantic image synthesis (Isola et al., 2017; Berrada et al., 2024) and super-resolution (Suvorov et al., 2022; Jo et al., 2020), and to assess the sample diversity of generative image models (Schönfeld et al., 2021; Astolfi et al., 2024). In addition, recent works propose variants that do not require pretrained image backbones (Amir & Weiss, 2021; Czolbe et al., 2020; Veeramacheneni et al., 2023). An et al. (2024); Song et al. (2023) employed LPIPS as metric function in pixel space to train cascaded diffusion models and consistency models, respectively. More closely related to our work, Kang et al. (2024) used a perceptual loss (E-LatentLPIPS) defined in latent space to distill LDMs to conditional GANs, but used a separate image classification network trained over latents rather than the autoencoder’s decoder to obtain the features for this loss. Lin & Yang (2024) added a perceptual loss (Self-Perceptual) to train LDMs that is defined over the features of the denoiser network in the case of UNet architectures. However, they found this loss to be detrimental when using classifier-free guidance for inference. In summary, compared to prior work on perceptual losses, our work is different in that (i) LPL is defined over the features of the decoder – that maps from latent space to RGB pixel space, rather than using a network that takes RGB images as input – and (ii) our LPL can be regardlessly applied to train both latent diffusion and flow models and has no architecture constraint in the denoiser model.
@@ 3 Using the latent decoder to define a perceptual loss
3.1 Latent diffusion and the $\ell_{2}$ objective
@@
We use $F_{\beta}$ to refer to an autoencoder that consists of two modules. The encoder, $F^{e}_{\beta}$, maps RGB images ${\bf x}_{0}\in\mathbb{R}^{H\times W\times 3}$ to a latent representation ${\bf z}_{0}\in\mathbb{R}^{H/d\times W/d\times C}$, where $d$ is the spatial downscaling factor, and $C$ the channel dimension of the autoencoder. The decoder, $F^{d}_{\beta}$, maps from the latent space to the RGB image space. In LDM, the diffusion model, $D_{\Theta}$, with parameters $\Theta$ is defined over the latent representation of the autoencoder. We follow a typical setting, see e.g. (Peebles & Xie, 2023; Chen et al., 2024; Rombach et al., 2022), where we use a fixed pre-trained autoencoder with a downsampling factor of $d=8$ and a channel capacity of $C=4$.
@@
Training Objective. The diffusion formulation results in an objective function that is a lower-bound on the log-likelihood of the data. In the DDPM paradigm (Ho et al., 2020), the variational lower bound can be expressed as a sum of denoising score matching losses (Vincent, 2011), and the objective function can be written as $\mathcal{L}=\sum_{t}\mathcal{L}_{t}$, where $\mathcal{L}_{t}=\mathcal{D}_{\textrm{KL}}\left[q\left({\bf x}_{t-1}|{\bf x}_{t},{\bf x}_{0}\right)||p_{\theta}\left({\bf x}_{t-1}|{\bf x}_{t}\right)\right]=\mathbb{E}_{{\bf x}_{0},\boldsymbol{\epsilon},t}\left[\frac{\beta_{t}}{{(1-\beta_{t})(1-\alpha_{t})}}\cdot\lVert{\boldsymbol{\epsilon}}-\epsilon_{\theta}({\bf x}_{t},t)\rVert^{2}\right]$. Ho et al. (2020) observed that disregarding the time step specific weighting resulted in improved sample quality, and introduced a simplified noise reconstruction objective, known as epsilon prediction, where the objective is the average of the MSE loss between the predicted noise and the noise vector added to the image, $\mathcal{L}_{\text{simple}}=\sum_{t}\mathbb{E}_{{\bf x}_{0},\boldsymbol{\epsilon},t}\left[\lVert\boldsymbol{\epsilon}-\epsilon_{\theta}({\bf x}_{t},t)\rVert^{2}\right]$. The underlying idea is that the better the noise estimation, the better the final sample quality. An equivalent way to interpret this objective is through reparameterization of the target to the original latents, $\mathcal{L}_{\text{simple}}=\sum_{t}\lambda_{t}\cdot\mathbb{E}_{{\bf x}_{0},\boldsymbol{\epsilon},t}\left[\lVert{\bf x}_{0}-\hat{x}_{0}({\bf x}_{t},t;\theta)\rVert^{2}\right]$, where $\hat{x}_{0}({\bf x}_{t},t;\theta)=({\bf x}_{t}-\sigma_{t}\hat{\boldsymbol{\epsilon}}_{t})/\alpha_{t}$ and $\lambda_{t}=1/\sigma_{t}^{2}$.
@@
We note that the presence of $\ell_{2}$ in the LDM objective has two important implications. First, the $\ell_{2}$ norm treats all pixels in the latents as equally important and disregards the downstream structure induced by the decoder whose objective is to reconstruct the image from its latents. This is problematic because the autoencoder’s latent space has a non-uniform structure and is not equally influenced by the different pixels in the latent code. Thus, optimizing the $\ell_{2}$ distance in the diffusion model latent space could be different from optimizing the perceptual distance between images. Second, while an $\ell_{2}$ objective is theoretically justified in the original DDPM formulation, generative models trained with an $\ell_{2}$ reconstruction objective have been observed to produce blurry images, as is the case e.g. for VAE models (Kingma & Welling, 2014). Such an effect can lead to exposure bias in diffusion (i.e. denoiser input mismatch between training and sampling), which has been studied through the lens of the sampling algorithm but not the training objective Ning et al. (2024). The problem of blurry images due to $\ell_{2}$ reconstruction losses has been addressed through the use of perceptual losses such as LPIPS (Zhang et al., 2018), which provide a significant boost to the image quality in settings such as autoencoding (Esser et al., 2021), super-resolution Ledig et al. (2017) and image-to-image generative models (Isola et al., 2017; Park et al., 2019).
To address these two implications of the latent $\ell_{2}$ optimization, we design a new loss, namely Latent Percetual Loss (LPL), that directly tackles the mismatch between the structures of the latent and the pixel space by matching the target and the predicted latents when partially decoded with the AE decoder. Compared to existing perceptual losses, e.g., LPIPS, that usually rely on features from pre-trained classifier networks, the use of AE decoder features is not standard. However, for LDMs training, the AE decoder provides a perceptually meaningful signal as the intermediate features of the decoder, being closer to pixel space, have a structure more similar to the pixel space. Moreover, these features have higher resolution than the diffusion latent features, which helps in modeling high-level details that reduce blurriness of the predicted image.
@@ 3.2 Latent perceptual loss
We propose a loss function that operates on the features at different depths in the autoencoder’s decoder, $F^{d}_{\beta}$. Let ${\bf z}_{t}=\alpha_{t}{\bf z}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t}$ be a noisy sample in the diffusion model latent space at time $t$, and $\hat{{\bf z}}_{0}=({\bf z}_{t}-\sigma_{t}D({\bf z}_{t},t;\Theta))/\alpha_{t}$ the corresponding estimated noise-free latent at time $t=0$.
To enhance the accuracy of the reconstruction process in image space, we propose augmenting the diffusion objective with a penalty term that encourages the predicted forward process distribution to match the reconstructions $\hat{{\bf z}}_{0}$ at a lower level of compression. This is achieved by introducing an image-level projection penalty on the forward process, which can be expressed as:
$\mathcal{L}_{t-1}^{\text{pen}}=\mathbb{E}_{q}\left[D_{\text{KL}}\left(q({{\bf x}}_{t-1}|{{\bf x}}_{t},{{\bf x}}_{0})\parallel p_{\Theta}({\hat{\bf x}}_{t-1}|{\hat{\bf x}}_{t})\right)\right].$ (1)
In Section A.1, we show that such a penalty, under certain conditions, can be approximated by computing a pairwise distance between the outputs of the projection decoder, which maps latents into samples in image-space. Guided by this result, our loss term should resemble a reconstruction term between the original image and the predicted/denoised image obtained from the decoder of the autoencoder. However, this does not address the problem of blurriness exposed in the previous section, which is why we opt for a loss structure that is similar to LPIPS, where the loss term is defined over the intermediate feature spaces of the decoder.
@@
We compute two sets of hierarchies of $L$ decoder features, $\{\boldsymbol{\phi}_{l}\}_{l=1}^{L}$, and, $\{\hat{\boldsymbol{\phi}}_{l}\}_{l=1}^{L}$, by decoding both the original, ${\bf z}_{0}$, and estimated latents, $\hat{{\bf z}}_{0}$ (where for brevity we drop the dependence of $\hat{{\bf z}}_{0}$ on $t$):
$\begin{cases}\boldsymbol{\phi}_{1},...,\boldsymbol{\phi}_{L}=\big{(}F_{\beta}^{d,l}({\bf z}_{0})\big{)}_{l\in[\![1,L]\!]},\\ \hat{\boldsymbol{\phi}}_{1},...,\hat{\boldsymbol{\phi}}_{L}=\big{(}F_{\beta}^{d,l}(\hat{{\bf z}}_{0})\big{)}_{l\in[\![1,L]\!]}.\end{cases}$ (2)
@@
$\mathcal{L}_{\textrm{LPL}}=\mathbb{E}_{t\in\mathcal{T},\boldsymbol{\epsilon}\sim\mathcal{N}(0,I),{\bf x}_{0}\in D_{\mathcal{X}}}\left[\delta_{\sigma_{t}\leq\tau_{\sigma}}\sum_{l=1}^{L}\frac{\omega_{l}}{C_{l}}\sum_{c=1}^{C_{l}}\big{\lVert}\rho_{l,c}(\hat{\phi}_{l,c})\odot\big{(}\boldsymbol{\phi}^{\prime}_{l,c}-\hat{\boldsymbol{\phi}}^{\prime}_{l,c}\big{)}\big{\rVert}_{2}^{2}\right].$ (3)
where $\boldsymbol{\phi}^{\prime}_{l}$ is the standardized version of $\boldsymbol{\phi}_{l}$ across the channel dimension, $\rho_{l,c}(\hat{\boldsymbol{\phi}}_{l,c})$ is a binary map masking the detected outliers in the feature map $\hat{\boldsymbol{\phi}}_{l,c}$, $\omega_{l}$ is a depth-specific weighting and $C_{l}$ the channel dimensionality of the feature tensor. Note that we better explain these terms later in this section. Moreover, to reduce both LPL computational complexity and memory overhead, we only apply our loss for high signal-to-noise ratios (SNR). In particular, we impose a hard threshold $\tau_{\sigma}$ and only apply the loss if the SNR is higher than it $\delta_{\sigma_{t}\leq\tau_{\sigma}}(\sigma_{t})$. The LPL loss is applied in conjunction with the standard diffusion loss, resulting in the training objective
$\mathcal{L}_{\textrm{tot}}=\mathcal{L}_{\textrm{Diff}}+w_{\textrm{LPL}}\cdot\mathcal{L}_{\textrm{LPL}}$.
@@
Framework $\alpha_{t}$ $\sigma_{t}$ $\hat{{\bf x}}_{0}$
DDPM-$\epsilon_{t}$ $\sqrt{\bar{\alpha}_{t}}$ $\sqrt{1-\bar{\alpha}_{t}}$ $\big{(}{\bf x}_{t}-\sigma_{t}D({\bf x}_{t},t;\Theta)\big{)}/\alpha_{t}$
DDPM-$v_{\theta}$ $\sqrt{\bar{\alpha}_{t}}$ $\sqrt{1-\bar{\alpha}_{t}}$ $\alpha_{t}{\bf x}_{t}-\sigma_{t}D({\bf x}_{t},t;\Theta)$
Flow-OT $1-t$ $t$ ${\bf x}_{t}-\sigma_{t}D({\bf x}_{t},t;\Theta)$
Table 1: Summary of the formula for the estimate of the clean image corresponding to the different formulations. Using the following parameterization, $\forall t,{\bf x}_{t}=\alpha_{t}{\bf x}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t}$.